Child heights. Heights of 10 year olds, regardless of gender, closely follow a normal distribution with mean 55 inches and standard deviation 7 inches.

(a) If the tallest 10% of the class is considered "very tall", what is the height cutoff for "very tall"? ___________ (Round to one decimal place)

(b) The height requirement for Batman the Ride at Six Flags Magic Mountain is 57 inches. What proportion of 10 year olds cannot go on this ride? In other words, what is probability a random 10 year old is below 57 inches? _______________(Use three significant figures.)
Answer #1

Facts : 1) If X and Z are the normal and standard normal distibuted random variables respectively. X-μ X-interval can be conva) Now we have to find the value of x such that P(X > x) = 10% = 0.1 P(X > x) = P(X – u > x-u) = P (**> *74) x- = P(Z > z) saUsing Excel formula: ENORM.S.INV(0.9), we got z = 1.28. =z= 1.28 x-u = 1.28 o X-55 = 1.28 7 = x= 7* (1.28) +55 x= 64.0 .: x=b) Now we have to find P(X<57). P(X<57) = P(X – u <57 – 55) = P < = P(Z < 0.29) 57-55 = 0.614 .: P(X<57) = 0.614

thank u